Solve for the unlisted blood pressure in each subject. INCLUDE YOUR UNITS FOR EACH VALUE

Subject           Systolic pressure                  Diastolic pressure                Pulse pressure

1                                  130                                          80                                ____(E)______

2                                  110                                          ____(F)____                           40

3                                  ____(G)____                           66                                            38

4                                  140                                          ____(H)____                           64
Answer #1

1. Pulse pressure = Systolic pressure - diastolic pressure

                             = 130 - 80

(E) is 50 mmHg

2. Diastolic pressure = Systolic pressure - pulse pressure

                                 = 110 - 40

(F) is 70 mm Hg

3. Systolic pressure = Diastolic pressure + pulse pressure

                                 = 66 + 38

(G) is 104 mm Hg

4. Diastolic pressure = systolic pressure - pulse pressure

                                  = 140 - 64

(H) is 76 mm Hg